One step in producing a spare part, involves drilling 4 holes. From 150 samples, the average time needed for this process is 72 seconds and SD = 10 seconds.
(a) calculate the CIs each with 95% CIs and 99.5% for the average time needed for the process.
(b) How many samples must be taken to get a 99.5% CI for averages in Ŧ 1.5 seconds from the actual price?
1
Expert's answer
2020-05-04T08:11:21-0400

The Central Limit Theorem

Let "X_1,X_2,...,X_n" be a random sample from a distribution with mean "\mu" and variance "\sigma^2." Then if "n"

is sufficiently large, "\bar{X}" has approximately a normal distribution with "\mu_{\bar{X}}=\mu" and "\sigma_{\bar{X}}^2=\sigma^2/n."

If "n>30" the Central Limit Theorem can be used.

The following information is provided: "\bar{X}=72, \sigma=10, n=150."

We need to construct the 95% confidence interval for the population mean "\mu."

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96." The corresponding confidence interval is computed as shown below:


"CI=\big(\bar{X}-z_c{\sigma\over\sqrt{n}},\bar{X}+z_c{\sigma\over\sqrt{n}}\big)="

"=\big(72-1.96\times{10\over\sqrt{150}},72+1.96\times{10\over\sqrt{150}}\big)\approx"

"\approx(70.400, 73.600)"

We need to construct the 99.5% confidence interval for the population mean "\mu."

The critical value for "\alpha=0.005" is "z_c=z_{1-\alpha/2}=2.807."

The corresponding confidence interval is computed as shown below:


"CI=\big(\bar{X}-z_c{\sigma\over\sqrt{n}},\bar{X}+z_c{\sigma\over\sqrt{n}}\big)="

"=\big(72-2.807\times{10\over\sqrt{150}},72+2.807\times{10\over\sqrt{150}}\big)\approx"

"\approx(69.708, 74.292)"

(b) How many samples must be taken to get a 99.5% CI for averages in "\pm1.5" seconds from the actual price?


"z_c{\sigma\over\sqrt{n}}\leq1.5"

"n\geq\bigg({z_c\sigma\over1.5}\bigg)^2"

"n\geq\bigg({2.807\cdot10\over1.5}\bigg)^2"

"n\geq351"